# Images Speak Louder than Words: Understanding and Mitigating Bias in Vision-Language Model from a Causal Mediation Perspective

### v2

@@ Abstract
Vision-language models (VLMs) pre-trained on extensive datasets can inadvertently learn biases by correlating gender information with specific objects or scenarios. Current methods, which focus on modifying inputs and monitoring changes in the model’s output probability scores, often struggle to comprehensively understand bias from the perspective of model components. We propose a framework that incorporates causal mediation analysis to measure and map the pathways of bias generation and propagation within VLMs. Our framework is applicable to a wide range of vision-language and multimodal tasks. In this work, we apply it to the object detection task and implement it on the GLIP model. This approach allows us to identify the direct effects of interventions on model bias and the indirect effects of interventions on bias mediated through different model components. Our results show that image features are the primary contributors to bias, with significantly higher impacts than text features, specifically accounting for 32.57% and 12.63% of the bias in the MSCOCO and PASCAL-SENTENCE datasets, respectively. Notably, the image encoder’s contribution surpasses that of the text encoder and the deep fusion encoder. Further experimentation confirms that contributions from both language and vision modalities are aligned and non-conflicting. Consequently, focusing on blurring gender representations within the image encoder which contributes most to the model bias, reduces bias efficiently by 22.03% and 9.04% in the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal performance loss or increased computational demands.
@@ 1 Introduction
Vision-language models have shown promising results in tasks such as classification Li et al. (2023); Jia et al. (2021); Radford et al. (2021), image search Sun et al. (2023); Radford et al. (2021); Li et al. (2023), and object detection Kuo et al. (2023); Li et al. (2022) by training on large-scale image-text pairs to understand the correspondences between cross-modal image features and language features. Models trained on extensive datasets exhibit excellent zero-shot capabilities Radford et al. (2021); Yu et al. (2022); Li et al. (2022); Zhang et al. (2022a) but also risk discovering and exploiting societal biases present in the underlying image-text pair corpora, potentially introducing bias that leads to social unfairness  Zhao et al. (2017). The revelation, measurement, and understanding of biases within models  Zhou et al. (2022); Zhang et al. (2022b); Lee et al. (2023); Vig et al. (2020) has sparked widespread research interest and are crucial for bias mitigation  Zhang et al. (2022b); Seth et al. (2023); Dehdashtian et al. (2023). However, most contemporary methods, derived from language models, lack standardized metrics for evaluating bias and primarily assess the correlation between the outputs of classifiers and external attributes  Zhang et al. (2022b).  Barrett et al. (2019) noted that interpretations based on classifier outputs can be factually inaccurate and not generalizable. While these methods can highlight the impacts of certain contributions on model outputs, they (1) fail to comprehend the generation and flow of bias within the model and (2) do not understand the causal roles of model components in the generation and propagation of bias. Consequently, they are not able to provide clear guidance on how to effectively mitigate bias at the model level.
@@
In this work, we propose a standardized framework to measure bias in VLMs, providing a comprehensive understanding of how bias flows within the entire model structure. Specifically, we use the GLIP model Li et al. (2022) as a case study, focusing on gender bias in the task of object detection, which is a predominant and challenging problem in computer vision. We conduct the analysis on both the MS-COCO Lin et al. (2014) and PASCAL-SENTENCE Rashtchian et al. (2010) datasets. We observe that GLIP model exhibits unbalanced inference capabilities on different genders, with certain indoor object categories like pets more likely to be associated with females and outdoor objects like vehicles with males. To holistically understand how the bias flows in the model, we adapt causal mediation analysis  Vig et al. (2020) to VLMs, providing a finer-grained study of contributions from different model components. We find that, among the different model components (text module, image module, and fusion module that combines them), the image module contributes the most to the model’s bias – over twice as much as the text module. In the MSCOCO and PASCAL-SENTENCE datasets, image features accounted for 32.57% and 12.63% of the bias generated, respectively, compared to approximately 15.48% and 5.64% by text features. Also, the interaction and updating process between image and text features during the deep fusion process significantly impacts bias production, accounting for about 57% of the contributions in the image and text encoders.
Furthermore, by integrating interventions across different modules, we discovered that their contributions to bias are aligned and do not conflict, allowing us to prioritize bias mitigation efforts within the image encoder, which is the most substantial contributor to bias. Based on the results, we propose to effectively mitigate the bias in VLMs: reducing biases from the image module can successfully reduce bias by 22.03% on the MSCOCO dataset and 9.04% on the PASCAL-SENTENCE dataset, compared to a reduction of 7.08% and 1.18% in the text module. In summary, the contributions of our work are:
@@ 2 Related Work
In recent years, vision-language models (VLMs) have experienced rapid advancements. The latest developments in VLMs often employ a dual-stream architecture that separately encodes text and images Kim et al. (2021), and these are then merged and aligned to facilitate cross-modal understanding of visual and linguistic features  Radford et al. (2021). Furthermore, some studies treat the joint training of images and text as a phrase localization process, aiming to better align and integrate visual and linguistic features  Li et al. (2022). Typically, these models are trained on image-text pairs from datasets such as MSCOCO  Lin et al. (2014), VQA  Antol et al. (2015), OpenImages  Kuznetsova et al. (2020), and Flickr30k Entities  Plummer et al. (2015), achieving impressive results in various downstream tasks including image classification  Radford et al. (2021), image generation  Radford et al. (2021), visual question answering  Li et al. (2018); Antol et al. (2015), and image captioning  Lu et al. (2019); Alayrac et al. (2022).
@@ 4 Experimental Setup of Bias Measurement and Understanding
Initially, we implement replace-gender and mask-gender interventions on the inputs respectively without any alterations to the model components. By monitoring the changes in the values of BiasVL, the individual impacts of image and text inputs on gender bias within the input module were assessed. Subsequently, we conducted a detailed causal mediation analysis on the text encoder and image encoder. Previous studies  Vaswani (2017); Vig and Belinkov (2019) have emphasized the critical roles that attention heads and layers play in deep learning models. Therefore, we selected attention heads from these two modules as mediators. Specifically, we chose the attention heads within a specific layer, along with those in all preceding layers as mediators, and conducted experiments from shallow to deep layers. This analysis aimed to identify whether the text encoder or image encoder contributes more significantly to gender bias and to determine which layers in the model are principally responsible for bias generation. It also sought to understand how bias flows and accumulates across different layers within the encoders. Then, we selected a combination of attention layers from both the image encoder and text encoder as mediators to observe changes in bias and compare these results with previous findings, exploring whether different modalities reinforce bias or conflict in the direction of bias.
@@ 5.1 Bias Measurement
Dataset BiasVL BiasVL(M1,M2) BiasVL(F1,F2) MSCOCO 1.434 0.278 0.359 PASCAL-S 1.369 0.341 0.381
@@ 5.2 Bias Understanding with Causal Mediation Analysis
Applying the mask-gender intervention to the input image module reduced the BiasVL to 0.967 for the MSCOCO dataset and to 0.664 for the PASCAL-SENTENCE dataset, representing reductions of approximately 32.57% and 12.63%, respectively. We employed the attention heads in the image encoder as the mediator
to examine both the indirect effects of this model component and the direct effects of the mask-gender on predictions. Figure 55(a) and Figure 55(e) illustrate that employing more attention heads as mediators leads to an increase in the value of the indirect effect, while the direct effect diminishes. This supports an intuition that removing gender information from more layers in the image encoder weakens the model’s dependency on latent correlations between gender in images and specific objects, thus mitigating gender bias in predictions. Furthermore, while interventions at the input level significantly impact final predictions, targeting the image encoder alone achieves about 53% of the mask-gender effect.
@@
To further validate whether image features contribute more to bias generation than text features, we utilized the attention heads in the deep fusion encoder as the mediator, adjusting the attention heads’ parameters in the states of either mask-gender intervention or replace-gender intervention. The results displayed in Figure 55(d) and Figure 55(c) show that for the MSCOCO dataset, the indirect effects from mask-gender and replace-gender through the deep fusion encoder are up to 0.260 and 0.189, respectively, reducing the BiasVL by approximately 18.13% and 13.18%. For the PASCAL-Sentence dataset, the reductions are 10.80% and 0.53%, respectively (Figure 55(h) and Figure 55(g)). These findings reaffirm our conclusion that image features play a more substantial role in bias generation than text features. They also suggest that even though the deep fusion module does not extract features directly from images and text, the interactive updating process between text and image features significantly influences bias generation, accounting for approximately 55.70% of the effect observed with the encoder alone.
@@
GLIP_ImageFair 46.2 17.3 1.118 0.694 22.03% 9.04%
GLIP_TextFair 46.6 17.6 1.322 0.754 7.8% 1.18%
@@ 8 Results
As indicated in Table 2, blurring gender representations in the image encoder demonstrated significant bias mitigation on both the MSCOCO and PASCAL-SENTENCE datasets. The experimental findings suggest that obscuring gender information in the image encoder is more effective at reducing model bias compared to similar interventions in the text encoder. Our results show that by blurring gender representations in the image features within the image encoder, we effectively reduced model bias by 22.03% and 9.04% on the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal impact on model performance.

### v3

@@ Abstract
Vision-language models (VLMs) pre-trained on extensive datasets can inadvertently learn biases by correlating gender information with specific objects or scenarios. Current methods, which focus on modifying inputs and monitoring changes in the model’s output probability scores, often struggle to comprehensively understand bias from the perspective of model components. We propose a framework that incorporates causal mediation analysis to measure and map the pathways of bias generation and propagation within VLMs. Our framework is applicable to a wide range of vision-language and multimodal tasks. In this work, we apply it to the object detection task and implement it on the GLIP model. This approach allows us to identify the direct effects of interventions on model bias and the indirect effects of interventions on bias mediated through different model components. Our results show that image features are the primary contributors to bias, with significantly higher impacts than text features, specifically accounting for approximately 33% and 13% of the bias in the MSCOCO and PASCAL-SENTENCE datasets, respectively. Notably, the image encoder’s contribution surpasses that of the text encoder and the deep fusion encoder. Further experimentation confirms that contributions from both language and vision modalities are aligned and non-conflicting. Consequently, focusing on blurring gender representations within the image encoder which contributes most to the model bias, reduces bias efficiently by approximately 22.03% and 9.04% in the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal performance loss or increased computational demands.
@@ 1 Introduction
Vision-language models have shown promising results in tasks such as classification Li et al. (2023); Jia et al. (2021); Radford et al. (2021), image search Radford et al. (2021); Li et al. (2023), and object detection Kuo et al. (2023); Li et al. (2022) by training on large-scale image-text pairs to understand the correspondences between cross-modal image features and language features. Models trained on extensive datasets exhibit excellent zero-shot capabilities Radford et al. (2021); Yu et al. (2022); Li et al. (2022); Zhang et al. (2022a) but also risk discovering and exploiting societal biases present in the underlying image-text pair corpora, potentially introducing bias that leads to social unfairness  Zhao et al. (2017). The revelation, measurement, and understanding of biases within models  Zhou et al. (2022); Zhang et al. (2022b); Lee et al. (2023); Vig et al. (2020) has sparked widespread research interest and are crucial for bias mitigation  Zhang et al. (2022b); Seth et al. (2023); Dehdashtian et al. (2024). However, most contemporary methods, derived from language models, lack standardized metrics for evaluating bias and primarily assess the correlation between the outputs of classifiers and external attributes  Zhang et al. (2022b).  Barrett et al. (2019) noted that interpretations based on classifier outputs can be factually inaccurate and not generalizable. While these methods can highlight the impacts of certain contributions on model outputs, they (1) fail to comprehend the generation and flow of bias within the model and (2) do not understand the causal roles of model components in the generation and propagation of bias. Consequently, they are not able to provide clear guidance on how to effectively mitigate bias at the model level.
@@
In this work, we propose a standardized framework to measure bias in VLMs, providing a comprehensive understanding of how bias flows within the entire model structure. Specifically, we use the GLIP model Li et al. (2022) as a case study, focusing on gender bias in the task of object detection, which is a predominant and challenging problem in computer vision. We conduct the analysis on both the MS-COCO Lin et al. (2014) and PASCAL-SENTENCE Rashtchian et al. (2010) datasets. We observe that GLIP model exhibits unbalanced inference capabilities on different genders, with certain indoor object categories like pets more likely to be associated with females and outdoor objects like vehicles with males. To holistically understand how the bias flows in the model, we adapt causal mediation analysis  Vig et al. (2020) to VLMs, providing a finer-grained study of contributions from different model components. We find that, among the different model components (text module, image module, and fusion module that combines them), the image module contributes the most to the model’s bias – over twice as much as the text module. In the MSCOCO and PASCAL-SENTENCE datasets, image features accounted for approximately 32.57% and 12.98% of the bias generated, respectively, compared to approximately 15.48% and 5.64% by text features. Also, the interaction and updating process between image and text features during the deep fusion process significantly impacts bias production, accounting for about 56% of the contributions in the image and text encoders. Furthermore, by integrating interventions across different modules, we discovered that their contributions to bias are aligned and do not conflict, allowing us to prioritize bias mitigation efforts within the image encoder, which is the most substantial contributor to bias. Based on the results, we propose to effectively mitigate the bias in VLMs: reducing biases from the image module can successfully reduce bias by approximately 22.03% on the MSCOCO dataset and 9.04% on the PASCAL-SENTENCE dataset, compared to a reduction of approximately 7.81% and 1.18% in the text module. In summary, the contributions of our work are:
@@ 2 Related Work
In recent years, vision-language models (VLMs) have experienced rapid advancements. The latest developments in VLMs often employ a dual-stream architecture that separately encodes text and images Kim et al. (2021), and these are then merged and aligned to facilitate cross-modal understanding of visual and linguistic features  Radford et al. (2021). Furthermore, some studies treat the joint training of images and text as a phrase localization process, aiming to better align and integrate visual and linguistic features  Li et al. (2022). Typically, these models are trained on image-text pairs from datasets such as MSCOCO  Lin et al. (2014), VQA  Antol et al. (2015), OpenImages  Kuznetsova et al. (2020), and Flickr30k Entities  Plummer et al. (2015), achieving impressive results in various downstream tasks including image classification  Radford et al. (2021), image generation, visual question answering  Li et al. (2018); Antol et al. (2015), and image captioning  Lu et al. (2019); Alayrac et al. (2022).
@@ 4 Experimental Setup of Bias Measurement and Understanding
Initially, we implement replace-gender and mask-gender interventions on the inputs respectively without any alterations to the model components. By monitoring the changes in the values of BiasVL, the individual impacts of image and text inputs on gender bias within the input module were assessed. Subsequently, we conducted a detailed causal mediation analysis on the text encoder and image encoder. Previous studies  Vaswani et al. (2017); Vig and Belinkov (2019) have emphasized the critical roles that attention heads and layers play in deep learning models. Therefore, we selected attention heads from these two modules as mediators. Specifically, we chose the attention heads within a specific layer, along with those in all preceding layers as mediators, and conducted experiments from shallow to deep layers. This analysis aimed to identify whether the text encoder or image encoder contributes more significantly to gender bias and to determine which layers in the model are principally responsible for bias generation. It also sought to understand how bias flows and accumulates across different layers within the encoders. Then, we selected a combination of attention layers from both the image encoder and text encoder as mediators to observe changes in bias and compare these results with previous findings, exploring whether different modalities reinforce bias or conflict in the direction of bias.
@@ 5.1 Bias Measurement
Dataset BiasVL BiasVL(M1,M2) BiasVL(F1,F2) MSCOCO 1.434 0.278 0.359 PASCAL-S 0.763 0.341 0.381
@@ 5.2 Bias Understanding with Causal Mediation Analysis
Applying the mask-gender intervention to the input image module reduced the BiasVL to 0.967 for the MSCOCO dataset and to 0.664 for the PASCAL-SENTENCE dataset, representing reductions of approximately 32.57% and 12.98%, respectively. We employed the attention heads in the image encoder as the mediator to examine both the indirect effects of this model component and the direct effects of the mask-gender on predictions. Figure 55(a) and Figure 55(e) illustrate that employing more attention heads as mediators leads to an increase in the value of the indirect effect, while the direct effect diminishes. This supports an intuition that removing gender information from more layers in the image encoder weakens the model’s dependency on latent correlations between gender in images and specific objects, thus mitigating gender bias in predictions. Furthermore, while interventions at the input level significantly impact final predictions, targeting the image encoder alone achieves about 53% of the mask-gender effect.
@@
To further validate whether image features contribute more to bias generation than text features, we utilized the attention heads in the deep fusion encoder as the mediator, adjusting the attention heads’ parameters in the states of either mask-gender intervention or replace-gender intervention. The results displayed in Figure 55(d) and Figure 55(c) show that for the MSCOCO dataset, the indirect effects from mask-gender and replace-gender through the deep fusion encoder are up to 0.260 and 0.189, respectively, reducing the BiasVL by approximately 18.13% and 13.18%. For the PASCAL-Sentence dataset, the reductions are 10.80% and 0.53%, respectively (Figure 55(h) and Figure 55(g)). These findings reaffirm our conclusion that image features play a more substantial role in bias generation than text features. They also suggest that even though the deep fusion module does not extract features directly from images and text, the interactive updating process between text and image features significantly influences bias generation, accounting for approximately 55.7% of the effect observed with the encoder alone.
@@
GLIP_ImageFair 46.2 17.3 1.118 0.694 22.036% 9.043%
GLIP_TextFair 46.6 17.6 1.322 0.754 7.810% 1.180%
@@ 8 Results
As indicated in Table 2, blurring gender representations in the image encoder demonstrated significant bias mitigation on both the MSCOCO and PASCAL-SENTENCE datasets. The experimental findings suggest that obscuring gender information in the image encoder is more effective at reducing model bias compared to similar interventions in the text encoder. Our results show that by blurring gender representations in the image features within the image encoder, we effectively reduced model bias by approximately 22.036% and 9.043% on the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal impact on model performance.